Company XYZ purchased equipment and gave a three-year note with maturity value of $12,006. The annual discount rate for the note was 14% discounted semi-annually. Then the note price should be $10,000. True or False
Answer #1

Semi-annual discount rate = 14%/2 = 7%

Number of semi-annual periods = 3*2 = 6

Value of note = 12,006*PVF(7%, 6 years)

= 12,006*0.6663

= $8,000.10

Hence, the answer is FALSE

Note price should be $8000